OpenStudy (anonymous):

A certain product is packed in cubic boxes that measure 2 feet on a side. If a rectangular shipping container is 6 feet long, 4 feet wide, and 8 feet high, how many of these cubic boxes will be able to fit inside it?

OpenStudy (anonymous):

The smaller boxes are cubic boxes with side lengths of 2, meaning that each cubic box has a volume of 8 ft^3 (Volume = length*width*height, so V = 2*2*2 = 8). Next, you'd compute the volume of the shipping container, but before computing that, it's important to note that the width, height, and length are all multiples of two (the side lengths of the smaller boxes), meaning that a perfect amount of smaller boxes will be able to fit without empty space or without theoretically crushing the boxes to make them fit :P Now just calculate the volume of the shipping container (V = 6*4*8 = 192 ft^3). Just divide the total volume of the shipping container by the volume taken up by each one of the boxes and you should get an integer. 192 / 8 = 24. 24 boxes of a "certain product" would fit inside that shipping container :)

OpenStudy (anonymous):

thank you =]